# Accessing the topological properties of human brain functional sub-circuits in Echo State Networks

### v1

@@ 5.2 Read-out functional networks’ dynamics properties
Here, we evaluate read-out FN’s dynamics through diffusivity and communicability perspectives, measuring the extent to which information flows through the FN topology Estrada & Hatano (2008). Noticeably, we found that the communicability score reaffirms DMN dominance, not only in the statics domain (e.g., betweenness) but also in the dynamics domain (e.g., communicability). Nonetheless, DMN dominance is significantly larger (five times larger than the runner-up: VIS). This result would further strengthen the anticipated best performance from DMN for 3 out of 4 tasks.
@@ 5.3 Topological fitness of a priori set of FNs on connectome-based ESN
To evaluate the topological fitness of an a priori set of FN, we measure Q with pre-set $\sigma$ (e.g., Yeo’s FNs Yeo et al. (2011); Clauset et al. (2004a)). $Q=\sum_{i,j}(a_{ij}-p_{ij})(\delta_{\sigma_{i,u},\sigma_{j,v}})$ where $i,j$ are node indices; $u,v$ are FN indices; and $\sigma=[n]^{k}$ is the $n$-dimensional partition vector whose entries are indices from 1 to k (the number of FNs in an FC) We performed cross-comparison between processed functional and structural connectome on modularity measures, performing community detection and comparing adjusted mutual info score with the original parcellations. We found that, despite the functional networks parcellation derived from functional connectomes, the modularity of the structural topology is higher than functional (Figure 12), while structural connectomes also perform better than functional. We thus hypothesize that Yeo’s networks derived from Lausanne parcellation Suarez et al. (2020) and Schaefer parcellation Tipnis et al. (2021) are not information-theoretically aligned, or that topological fitness strongly determines model performance.
@@
6 Discussion and Conclusions
@@
In this study, we explored bio-plausible topology as computational structures in reservoir neural networks, particularly ESNs, and examined whether complex functional topologies contribute to performance in recurrent network settings. We found that different topological properties are conducive to the processing of different tasks. The topology of default mode network (DMN) facilitates better processing and memorization, implying a link to the ”functional dichotomy” of DMN found in traditional neuroscience studies Uddin et al. (2009). Functional networks are embedded and extensively tested in reservoir echo-state networks where different performance scaling and dynamic entry points are evaluated. In general, bigger connectomes perform better, albeit with diminishing returns, and the flow through functional networks determines how well the model performs. Experiments have also shown that the model performance is, to some extent, dependent on the absolute size of the readout functional networks. Further experiments on integrating functional signals to embedded reservoirs will be conducted to optimize the use of rs-fMRI or task fMRI data, more closely observing model performance and computational efficiency and expanding the model representation space using different activations.
Appendix A Additional Data Description
A.1 Reservoir Connectomic
For structural connectomes, we used connectomes from Suarez et al. (2020), publicly available at https://zenodo.org/records/4776453, which was originally processed from the University of Lausanne’s data, also available at: https://zenodo.org/records/2872624. For functional connectomes, we obtained the parcellated connectomes from the UCSD public repository: https://library.ucsd.edu/dc/object/bb59818382. All data are available for public use with proper citations (included in the main text). None of the authors of this paper overlapped with the authors of the published datasets mentioned.
A.2 COVID-19 Data
The ground truth for number of confirmed COVID-19 cases per region is gathered for four regions through open data:
We directly used the preprocessed final version of the data from Panagopoulos et al. (2021) GitHub repository in each country’s subfolder in the data folder, publicly available at: https://github.com/geopanag/pandemic_tgnn/tree/master/data.
Appendix B Detailed Experiment Settings
B.1 Framework
All of the code provided is in Python, with primary experiments run in conn2res (Suarez et al. (2023)) framework for reservoir computing. The specific implementation of conn2res that we used is publicly available at https://github.com/netneurolab/conn2res. Detailed implementation is included in the supplementary material as .zip file.
B.2 Hardware
All experiments were conducted on a single-CPU server with no GPU, with the following specifications:
  • CPU: 1x Ryzen 9 7950x
  • RAM: 64GB DDR5 4800MT/s
B.3 Hyperparameters
We provide the hyperparameters for our experiments w.r.t. the conn2res (Suarez et al. (2023)) framework. Our attached supplemental code also contains all original hyperparameters used in experiments.
  • Number of experiment runs $N=1000$
  • Reservoir input factor $C=0.001$
  • Train-test split: 80:20
  • Number of NeuroGymTask trials: $1000$ (PDM, PDMDR, CDM)
  • Number of MemCap trials: $4050$
  • Number of reservoir nodes (structural): 463 (450 cortical nodes + 13 subcortical nodes)
  • Ridge regressor hyperparameters: alpha=0.5, fit_intercept=False
Hyperparameters for traditional RNNs models are identical to main experiments. For full details on hyperparameters used, refer to attached supplemental code.
B.4 COVID-19 benchmark models
The models tested follow the same notation described in prior work on COVID-19 prediction (Panagopoulos et al. (2021)). The additional models we tested are:
Appendix C Network Measures
C.1 Betweenness centrality
Betweenness centrality is a measure of information routing through a network, where the shortest-path betweenness centrality is computed for nodes (Freeman (1977)). Betweenness centrality $c_{B}$ of a node $v$ is the sum of the fraction of all-pairs shortest paths that pass through $v$ (Brandes (2001)):
$c_{B}(v)=\sum_{s,t\in V}\frac{\sigma(s,t|v)}{\sigma(s,t)}$
where $V$ is the set of all nodes, $\sigma(s,t)$ is the number of shortest $(s,t)$-paths, and $\sigma(s,t|v)$ is the number of those paths passing through some node $v$ other than $s,t$. If $s=t$, $\sigma(s,t)=1$; if $v\in s,t$, $\sigma(s,t|v)=0$.
C.2 Communicability
Communicability is a measure modeling the extent of information transfer between nodes in the network, effectively provides a way to quantify the extent to which two regions in a network can communicate with each other (Estrada & Hatano (2008)). Communicability is computed using a spectral decomposition of the adjacency matrix. The communicability $C$ between two nodes $u$ and $v$ is computed using the connection between powers of the adjacency matrix and the number of walks:
$C(u,v)=\sum^{n}_{j=1}\phi_{j}(u)\phi_{j}(v)e^{\lambda_{j}}$
where $\phi_{j}(u)$ is the $u^{\text{th}}$ element of the $j^{\text{th}}$ orthonormal eigenvector of the adjacency matrix with eigenvalue $\lambda_{j}$
C.3 Modularity
Modularity is a measure of the degree to which the network can be separated into clearly independent groups. As previously defined and reduced (Clauset et al. (2004b)), the measure is defined as:
$Q=\sum^{n}_{c=1}\left[\frac{L_{c}}{m}-\gamma\left(\frac{k_{c}}{2m}\right)^{2}\right]$
where we iterate over all $c$ communities. $m$ is the numbed of network edges, $L_{c}$ is the number of within-community links for specific community $c$, $k_{c}$ is the sum of degrees of the nodes in community $c$, and $\gamma$ is the resolution parameter.
Appendix D Additional results
D.1 Experimental setup with functional connectomes
The resting state fMRI data from a single subject from the HCP1200 dataset pre-parcellated using the Kong2022 individual areal-level parcellation procedure were used in this study (Kong et al. (2021)). Models were examined under several parcellation settings, where the cortex is represented as functional connectomes resolutions ranging from 100 to 1000 nodes in increments of 100 nodes each (Kong et al. (2021)) and were evaluated on tasks. FC and adjacency matrix connections on each resolution describe the Pearson correlation between BOLD time series among brain regions. Regional labels and indices were collected from atlas mapping of the Schaefer atlas directory (Schaefer et al. (2017)), segregating every connectome resolution available of Kong’s parcellation into 17 functional Yeo functional networks (Thomas Yeo et al. (2011)): Visual (A, B, and C), Auditory, Somatomotor (SomMotA and B), Language, Salience/Ventral Attention (SalVenAttn A and B), Control (ContA, B, and C), Default (A, B, and C), and Dorsal Attention (DorsAttnA and B).
For a weighted rs-fMRI graph we construct several thresholded graphs on different thresholding values r ranging from -0.5 to 0.4 in increments of 0.1, with a threshold value of 0 effectively keeping only positively correlated connections on the original graph. Varying r effectively filters the graph, sparsifying graph connections based on the degree of thresholding in large degrees (magnitudes of connections filtered between two consecutive thresholds). Keeping negative correlations on lower threshold values may dilute the computation graph, and potentially decrease the model’s performance on several metrics. Additionally, each connectome is scaled by an alpha parameter between 0.05 and 2, which determines the value of the largest eigenvector of the connectomic adjacency matrix. The alpha scaling is applied after the normalization of all connections in the graph, and the effect of the alpha parameter on reservoir network dynamics is discussed in-depth in Suarez et al. (2020).
All models are evaluated on a perceptual decision-making task with delayed responsive, measuring both the computation and memory capacity of the reservoir networks (Inagaki et al. (2019)).
D.2 Reservoir size
In scaling experiments, the model generally performs better when larger parcellation sizes are used. Visualizations and numerical results indicated diminishing returns over later larger parcellations, while there is a substantial difference between the performance of the best and the worst alpha scaling values. Size efficiency decreases as resolution increases with diminishing returns starting at 400 nodes parcellation (Figure 18). On the other hand, SC reservoir with 1000 nodes has high performance at 0.82 F1-score while FC at 700 nodes already reaches 0.75 F1-score, demonstrating FC’s parameter efficiency and signifying that FCs can be better scaled to finer resolutions with higher model performance.
D.3 Functional thresholding
Hyperparameter combination experiments showed that functional connectome benefits greatly when positively thresholded, possibly confirming the hypothesis that negative correlations dampen reservoir dynamics. Reservoirs performed optimally when thresholded at 0-0.2 and scaled at alpha 0.85-1.15 (Figure LABEL:fig:hyperparams).
D.4 Permutations
Input-output permutation experiments significantly varied performance on different pairings. There are several different configurations where the model performed optimally for different parcellations; however, better performance depends on the set of output nodes with Control and Default outputs outperforming other groups. Most notable are readouts from Auditory and Language functional networks where performance is significantly worse than other readouts. Said two functional network groups have substantially fewer nodes than the other functional network groups, possibly indicating that performance heavily depends on the size of the readouts up to the optimal signal-to-noise ratio (Figure LABEL:fig:permutations).
D.5 Readout size
Readout balancing and permutation experiments show a potential correlation between the readout functional network size and the model’s performance. While equal readout functional networks perform roughly uniformly, the augmented functional network readouts perform slightly better with increasing size relative to other readout functional networks. Such is the case for the Language readouts from functional connectome and DA readouts from structural connectome, both of which did not perform well on the original connectomes. However, despite the general increase in performance and only slight variance between functional network readouts, certain combinations or networks performed worse than others regardless of readout partitioning. The performance anomaly refers to the darker 2-3 performance spots in the functional connectome, and similarly in structural connectome.
Figure 13: Structural 500 nodes, statistic-performance correlations, PDM, $\alpha=0.95$.
Figure 14: Structural 500 nodes, statistic-performance correlations, CDM, $\alpha=0.95$.
Figure 15: Structural 500 nodes, statistic-performance correlations, CDM, $\alpha=1.2$.
Figure 16: Structural 500 nodes, statistic-performance correlations, MemCap, $\alpha=0.95$.
Figure 17: Modularity histogram: partition algorithm greedy modularity or Louvain modularity-based, 20 bins, with adjusted mutual information w.r.t. the original partition.
Figure 18: Performance on tanh activation on the best (left) and worst (right) alpha values.

### v2

@@ 5.2 Read-out functional networks’ dynamics properties
Here, we evaluate read-out FN’s dynamics through diffusivity and communicability perspectives, measuring the extent to which information flows through the FN topology [56]. Noticeably, we found that the communicability score reaffirms DMN dominance, not only in the statics domain (e.g., betweenness) but also in the dynamics domain (e.g., communicability). Nonetheless, DMN dominance is significantly larger (five times larger than the runner-up: VIS). This result would further strengthen the anticipated best performance from DMN for 3 out of 4 tasks.
@@ 5.3 Topological fitness of a priori set of FNs on connectome-based ESN
To evaluate the topological fitness of an a priori set of FN, we measure Q with pre-set $\sigma$ (e.g., Yeo’s FNs [5, 59]). $Q=\sum_{i,j}(a_{ij}-p_{ij})(\delta_{\sigma_{i,u},\sigma_{j,v}})$ where $i,j$ are node indices; $u,v$ are FN indices; and $\sigma=[n]^{k}$ is the $n$ dimensional partition vector whose entries are indices from 1 to k (the number of FNs in an FC).
@@
We performed cross-comparison between processed functional and structural connectome on modularity measures, performing community detection and comparing adjusted mutual info score with the original parcellations. We found that despite the parcellation of functional networks derived from functional connectomes, the modularity of the structural topology is higher than functional (Figure 12), while structural connectomes also perform better than functional. We thus hypothesize that Yeo’s networks derived from Lausanne parcellation [2] and Schaefer parcellation [49] are not information-theoretically aligned or that topological fitness strongly determines model performance.
6 Discussion and Conclusion
@@
In this study, we explored bio-plausible topology as computational structures in reservoir neural networks, particularly ESNs, and examined whether complex functional topologies contribute to performance in recurrent network settings. We found that different topological properties are conducive to the processing of different tasks. The topology of default mode network (DMN) facilitates better processing and memorization, implying a link to the "functional dichotomy" of DMN found in traditional neuroscience studies [60].
Functional networks are embedded and extensively tested in reservoir echo-state networks where different performance scaling and dynamic entry points are evaluated. In general, bigger connectomes perform better, albeit with diminishing returns, and the flow through functional networks determines how well the model performs. Experiments have also shown that the model performance is, to some extent, dependent on the absolute size of the readout functional networks. Further experiments on integrating functional signals to embedded reservoirs will be conducted to optimize the use of rs-fMRI or task fMRI data, more closely observing model performance and computational efficiency and expanding the model representation space using different activations. Future studies should address connectomics-induced ESNs with a wider variety of benchmark models and applications such as predicting neurodegenerative disease (e.g., Alzheimer’s disease) pathology [61], or genetics phenotype [62].